A proton accelerates from rest in a uniform electric field of 610 N/C. At some later time, its speed is 1.4 106 m/s.
(a) Find the magnitude of the acceleration of the proton.
m/s2

(b) How long does it take the proton to reach this speed?
µs

(c) How far has it moved in that interval?
m

(d) What is its kinetic energy at the later time?
J
Answer #1
Concepts and reason

The concept required to solve the given problem is electric field.

For the first part, compute the net force acting on the proton and then rearrange the expression to obtain the acceleration.

For the second part and third part use kinematic equations to determine the time and distance travelled by the proton.

For the last part, calculate the kinetic energy by taking a product of the mass and square of the velocity.

Fundamentals

Electric field: It is the region around a charge where its influence can be felt.

It is given by,

E=keqr2E = \frac{{{k_e}q}}{{{r^2}}}

Here, ke{k_e} is the coulomb’s constant, qq is the charge and rr is the distance of the point from the charge where the electric field needs to be calculated.

The force acting on a charge of magnitude qq in an electric field EE is given by,

F=qEF = qE

Equations of motion: The three equation of motion are:

1.v=v0+atv = {v_0} + at

2.y=v0t+12at2y = {v_0}t + \frac{1}{2}a{t^2}

3.v2u2=2ay{v^2} - {u^2} = 2ay

Here, vv is the final velocity, v0{v_0} is the initial velocity, aa is the acceleration due to gravity, yy is the distance travelled and tt is the time taken.

(a)

The net force acting on the proton released from rest will be given by,

Fnet=qE{F_{net}} = qE

Here, qq is the charge and EE is the electric field.

The force is also given by,

F=maF = ma

Here, mm is the mass and aa is the acceleration.

Substitute mama for FF in equation (1).

ma=qEa=qEm\begin{array}{c}\\ma = qE\\\\a = \frac{{qE}}{m}\\\end{array}

Substitute 1.6×1019C1.6 \times {10^{ - 19}}{\rm{ C}} forqq, 1.67×1027kg1.67 \times {10^{ - 27}}{\rm{ kg}}for mm and 610N/C610{\rm{ N / C}} for EE in the above equation.

a=(1.6×1019C)(610N/C)(1.67×1027kg)=5.84×1010m/s2\begin{array}{c}\\a = \frac{{\left( {1.6 \times {{10}^{ - 19}}{\rm{ C}}} \right)\left( {610{\rm{ N / C}}} \right)}}{{\left( {1.67 \times {{10}^{ - 27}}{\rm{ kg}}} \right)}}\\\\ = 5.84 \times {10^{10}}{\rm{ m / }}{{\rm{s}}^2}\\\end{array}

(b)

The time taken by the proton to reach the speed 1.4×106m/s1.4 \times {10^6}{\rm{ m / s}} is given by,

v=u+att=vua\begin{array}{l}\\v = u + at\\\\t = \frac{{v - u}}{a}\\\end{array}

Here, vv is the final speed, uu is the initial speed and aa is the acceleration.

Since the proton starts from rest, the initial speed will be, u=0m/su = 0{\rm{ m / s}}.

The final speed will be, v=1.4×106m/sv = 1.4 \times {10^6}{\rm{ m / s}}.

The acceleration will be,

a=5.84×1010m/s2a = 5.84 \times {10^{10}}{\rm{ m / }}{{\rm{s}}^2}

Substitute 5.84×1010m/s25.84 \times {10^{10}}{\rm{ m / }}{{\rm{s}}^2} for aa, 1.4×106m/s1.4 \times {10^6}{\rm{ m / s}} for vvand 0m/s0{\rm{ m / s}} for uu in the above equation.

t=(1.4×106m/s)(0m/s)(5.84×1010m/s2)=2.40×105s\begin{array}{c}\\t = \frac{{\left( {1.4 \times {{10}^6}{\rm{ m / s}}} \right) - \left( {0{\rm{ m / s}}} \right)}}{{\left( {5.84 \times {{10}^{10}}{\rm{ m / }}{{\rm{s}}^2}} \right)}}\\\\ = 2.40 \times {10^{ - 5}}{\rm{ s}}\\\end{array}

(c)

The distance moved by proton in the calculated time interval will be,

s=ut+12at2s = ut + \frac{1}{2}a{t^2}

Here, uu is the initial speed, aa is the acceleration and tt is the time.

Substitute 5.84×1010m/s25.84 \times {10^{10}}{\rm{ m / }}{{\rm{s}}^2} for aa, 2.40×105s2.40 \times {10^{ - 5}}{\rm{ s}} for ttand 0m/s0{\rm{ m / s}} for uu in the above equation.

s=(0m/s)(2.40×105s)+12(5.84×1010m/s2)(2.40×105s)2=16.8m\begin{array}{c}\\s = \left( {0{\rm{ m / s}}} \right)\left( {2.40 \times {{10}^{ - 5}}{\rm{ s}}} \right) + \frac{1}{2}\left( {5.84 \times {{10}^{10}}{\rm{ m / }}{{\rm{s}}^2}} \right){\left( {2.40 \times {{10}^{ - 5}}{\rm{ s}}} \right)^2}\\\\ = 16.8{\rm{ m}}\\\end{array}

(d)

The kinetic energy at the later time will be,

K=12mv2K = \frac{1}{2}m{v^2}

Here, mm is the mass and vv is the velocity.

Substitute 1.4×106m/s1.4 \times {10^6}{\rm{ m / s}} for vv and 1.67×1027kg1.67 \times {10^{ - 27}}{\rm{ kg}} for mm in the above equation.

K=12(1.67×1027kg)(1.4×106m/s)2=1.64×1015J\begin{array}{c}\\K = \frac{1}{2}\left( {1.67 \times {{10}^{ - 27}}{\rm{ kg}}} \right){\left( {1.4 \times {{10}^6}{\rm{ m / s}}} \right)^2}\\\\ = 1.64 \times {10^{ - 15}}{\rm{ J}}\\\end{array}

Ans: Part a

The magnitude of acceleration of the proton is5.84×1010m/s25.84 \times {10^{10}}{\rm{ m / }}{{\rm{s}}^2}.